# Learning Action-based Representations Using Invariance

### v1

@@ Abstract
Robust reinforcement learning agents using high-dimensional observations must be able to identify relevant state features amidst many exogeneous distractors. A representation that captures controllability identifies these state elements by determining what affects agent control. While methods such as inverse dynamics and mutual information capture controllability for a limited number of timesteps, capturing long-horizon elements remains a challenging problem. Myopic controllability can capture the moment right before an agent crashes into a wall, but not the control-relevance of the wall while the agent is still some distance away. To address this we introduce action-bisimulation encoding, a method inspired by the bisimulation invariance pseudometric, that extends single-step controllability with a recursive invariance constraint. By doing this, action-bisimulation learns a multi-step controllability metric that smoothly discounts distant state features that are relevant for control. We demonstrate that action-bisimulation pretraining on reward-free, uniformly random data improves sample efficiency in several environments, including a photorealistic 3D simulation domain, Habitat. Additionally, we provide theoretical analysis and qualitative results demonstrating the information captured by action-bisimulation.
@@ 2 Related Works
Action-based Representations. RL methods have directly leveraged action-relevant representations in several ways. This includes contingency awareness (Bellemare et al., 2012; Choi et al., 2018; Chuck et al., 2020; 2023), which is closely related to action controllability (Zhong et al., 2020) and control information measures like empowerment (channel capacity between actions and state) (Jung et al., 2011; Mohamed & Jimenez Rezende, 2015; Levy et al., 2023) or affordances (Cruz et al., 2016; Khetarpal et al., 2020; Nagarajan et al., 2020). Multi-step inverse models are most similar to action-bisimulation, but common multi-step inverse methods (Lamb et al., 2022; Islam et al., 2022; Koul et al., 2023) require selecting a specific $k$ for the multi-step horizon, potentially leaving critical control information on the table. Furhter, it has been shown that multi-step inverse models can be insufficient when the dynamics are periodic (Levine et al., 2024). Action-bisimulation uses a soft invariance metric to extend single-step models, which better preserves long-term controllability.
Bisimulation methods. Bisimulation describes future invariant state representations, originally applied to stationary representations (Larsen & Skou, 1989; Dean et al., 1997; Ferns et al., 2004), before being extended to continuous state MDPs (Ferns et al., 2011). Reward-based bisimulation methods have gained popularity through learned deep representations (Zhang et al., 2020b). This has been extended to non-optimal policies (Castro et al., 2021), with generalized value function bounds (Kemertas & Aumentado-Armstrong, 2021) and augmented with state discretization (Kemertas & Jepson, 2022) and clustering (Liu et al., 2023). Bisimulation-based methods have also been applied in different contexts: expert policy similarity (Agarwal et al., 2021; Bertran et al., 2022; Mazoure et al., 2021), goal-conditioned RL (Hansen-Estruch et al., 2022) and reward-action policy equivalence (Liao et al., 2023; Castro, 2020). While this work draws on reward-bisimulation, action-bisimulation is fundamentally offline and task-agnostic because it takes an expectation over actions, removing its dependance on any policy.
@@ 3 Preliminaries
A Markov decision process is defined by the tuple $\mathcal{M}\coloneqq(\mathcal{S},\mathcal{A},p,R)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space and $s\in\mathcal{S},a\in\mathcal{A}$ are states and actions respectively. $p(s^{\prime}|s,a)$ is the transition function which gives the probability of the next state $s^{\prime}$ given the current state and action $(s,a)$. The reward function $R(s,a)$ maps state and action to a scalar reward. A policy $\pi(a|s)$ is the probability of an action given the current state.
This work utilizes the following two phase paradigm: in the first phase, the agent first takes actions without access to the reward function $R(s,a)$ to generate a dataset of ordered state action tuples $\mathcal{D}\coloneqq\{(s^{(0)},a^{(0)}),\ldots(s^{(|\mathcal{D}|-1)},a^{(|\mathcal{D}|-1)})\}$. Then, a representation $\phi:\mathcal{S}\rightarrow\mathcal{Z}$ is learned from $\mathcal{S}$. In the second phase, the agent learns from extrinsic reward utilizing the learned representation.
@@
$d_{\text{r-bisim}}(s_{i},s_{j})=\max_{a}\underbrace{(1-c)\cdot|R(s_{i},a)-R(s_{j},a)|}_{\text{base case}}+\underbrace{cW_{1}(d)(p(\cdot|s_{i},a),p(\cdot|s_{j},a))}_{\text{recursive step}}.$ (3)
We propose a novel controllability-based relation, which replaces reward equivalence with single-step control equivalence. By replacing rewards in the equivalence, the relation is task-agnostic.
@@
$d_{\text{a-bisim}}(s_{i},s_{j},\psi)=\underbrace{(1-c)\cdot\|\psi_{(}s_{i})-\psi_{(}s_{j})\|_{1}}_{\text{base case}}+\underbrace{c\cdot\mathbb{E}_{a\sim U(\mathcal{A})}\left[W_{1}(p(\cdot|s_{i},a),p(\cdot|s_{j},a))\right]}_{\text{recursive step}}$ (6)
@@ 4.2 Action-Bisimulation Metric
In the off-policy setting, we propose using one of two expectations for the recursive step: over the uniform distribution of actions $E_{a\sim U(\mathcal{A})}$ or over the behavior distribution: $E_{a\sim\pi_{b}(s_{i})}$. The use of the behavioral distribution is in settings where random actions might overly restrict the distribution of observed states. In practice, these are computed using the empirical mean. Then, the action-controllability bisimulation metric using learned models is:
$d_{\text{a-bisim}}(s_{i},s_{j},\psi_{\theta},\phi_{\eta})=(1-c)\cdot\|\psi_{\theta}(s_{i})-\psi_{\theta}(s_{j})\|_{1}+c\cdot\mathbb{E}_{a\sim U(\mathcal{A})}\left[W_{1}(f(\phi_{\nu}(s_{i}),a),f(\phi_{\nu}(s_{j}),a))\right]$ (9)
@@
$L(\mathcal{D})=\frac{1}{N}\sum_{s_{i},s_{j}\sim\mathcal{D}}\|\phi_{\eta}(s_{i})-\phi_{\eta}(s_{j})\|_{1}-d_{\text{a-bisim}}(s_{i},s_{j},\psi,\phi).$ (10)
@@
  Input: Dataset without reward $(s,a,s^{\prime})\sim D$, initial encoder $\phi^{\bar{\theta}}(s)$
@@ 5 Experiments
We evaluate experiments in three domains illustrated in Figure LABEL:fig:downstream_rl. Nav2D is a 15x15 grid environment where the agent navigates using cardinal directions to the center of the grid, avoiding randomly generated 2x2 obstacles. Pointmaze (Fu et al., 2020) is a 2D Mujoco control environment where the agent takes actions to reach a goal location while being impeded by obstacles. We also investigate Distractor Pointmaze where the background in Pointmaze has been replaced with photorealisic distractions in the form of video clips. Finally, Habitat (Savva et al., 2019b) is a complex 3D environment where the agent must navigate through scans of human environments to reach a goal location. Additional environment details are in Appendix H (number of obstacles, goal/grid size, randomization, etc.).
@@ 5.2 Downstream Learning
As we can see in Figure LABEL:fig:downstream_rl, learning with the action-bisimulation representation outperforms other methods in terms of sample efficiency, even k-step controllability (ACRO), by a substantial margin. This provides evidence for hypotheses 1 and 2, that action-bisimulation learns useful representations which compares well with other methods. That reconstruction and data augmentation-based methods bVAE and CURL perform poorly is not unexpected: in this domain the agent is often small, so these methods achieve low reconstruction loss even when they omit the most important element: the agent position. On the other hand, SSI captures the agent position but is highly myopic, limiting transfer to downstream tasks. We hypothesize ACRO struggles because it relies on predicting action from two states separated by $k$ timesteps, which is ill-posed, especially when using a dataset of random actions. Additional details on baselines can be found in Appendix G.
@@ 5.3 Background Distractors
In this section we evaluate hypothesis 3: whether the action-bisimulation encoding is robust to distractors. We assess this through a modified Pointmass environment with a photorealistic visual background. The foreground, that is the agent, goal position, and obstacles, remain the same. We visualize the distractor environments in Figure LABEL:paired_images, where the agent has been exagerrated.
@@ 6 Conclusion
Controllability-capturing encodings for reinforcement learning are a promising direction for representation pretraining since they can be learned without reward but are still able to filter out uncontrollable distractors. However, existing methods either only capture short-term controllability or are dependent on demonstration data, which has implicit task bias. We introduce the action-bisimulation encoding, which builds off of myopic representations by enforcing recursive invariance to learn a supervision-free multi-step controllability representation. The empirical results in this work demonstrate how these encodings can be used to improve the sample efficiency, especially in domains with significant background distractors. The primary limitation of this method is the inverse dyanmics single-step model, which might not capture all controllable features, just a subset. This can result in the representation being agnostic to important task elements. Nonetheless, action-bisimulation is a novel invariance relation for capturing controllability.
@@ Appendix E Additional RL Results
As we can see, this provides evidence that action-bisimulation encodings are not limited to discrete actions.
@@
Appendix F Limitations of Inverse Dynamics Action Bisimulation
@@
For example, consider the scenario where in the top corner of the screen there is a small display of the last action that the agent takes. In this scenario, the inverse dynamics model is likely to learn to only pay attention to this part of the screen, ignoring other components such as the state of the agent. This is because paying attention to this part is a sufficient representation of actions, even though it does not capture all action-relevant information. Information-based methods such as Equation 16 are on possible solution for this, but we have found empirically that they do not seem to learn representations useful for RL.
Finally, while multi-step controllability is a powerful property, not all tasks require this kind of lookahead, and it is not clear that multi-step pretraining would outperform single-step or other baselines in these cases. For example, in the popular mujoco locomotion domains (Todorov et al., 2012), knowing about future control can often be distracting to the agent—all the relevant information is captured by determining how the current action will affect future actions. Domains where long-term control are useful, such as manipulation, can also be challenging for the current form of action-bisimulation because of the minimal sufficient information property of the single-step losses. Future work is aimed at investigating this in greater detail.
Action bisimulation captures a partial or minimal representation of the single-step controllability elements, but not the entirety. This is even the case with mutual information objectives, though the difference in that case is less pronounced. As a result, future work is for investigating action-controllable components that are not necessary for inverse dynamics (action prediction).
@@ I.3 Habitat
Reinforcement Learning Parameters
Parameter Value Algorithm PPO Clip 0.2 Entropy Coef. 0.01 Batch Size $256$ Learning Rate $0.00025$

### v3

@@ Abstract
Robust reinforcement learning agents using high-dimensional observations must be able to identify relevant state features amidst many exogeneous distractors. A representation that captures controllability identifies these state elements by determining what affects agent control. While methods such as inverse dynamics and mutual information capture controllability for a limited number of timesteps, capturing long-horizon elements remains a challenging problem. Myopic controllability can capture the moment right before an agent crashes into a wall, but not the control-relevance of the wall while the agent is still some distance away. To address this we introduce action-bisimulation encoding, a method inspired by the bisimulation invariance pseudometric, that extends single-step controllability with a recursive invariance constraint. By doing this, action-bisimulation learns a multi-step controllability metric that smoothly discounts distant state features that are relevant for control. We demonstrate that action-bisimulation pretraining on reward-free, uniformly random data improves sample efficiency in several environments, including a photorealistic 3D simulation domain, Habitat. Additionally, we provide theoretical analysis and qualitative results demonstrating the information captured by action-bisimulation. Code and video: https://maxrudolph1.github.io/action-bisimulation-site/
@@ 2 Related Works
Action-based Representations. RL methods have directly leveraged action-relevant representations in several ways. This includes contingency awareness (Bellemare et al., 2012; Choi et al., 2018; Chuck et al., 2020; 2023), which is closely related to action controllability (Zhong et al., 2020) and control information measures like empowerment (channel capacity between actions and state) (Jung et al., 2011; Mohamed & Jimenez Rezende, 2015; Levy et al., 2023) or affordances (Cruz et al., 2016; Khetarpal et al., 2020; Nagarajan et al., 2020). Multi-step inverse models are most similar to action-bisimulation, but common multi-step inverse methods (Lamb et al., 2022; Islam et al., 2022; Koul et al., 2023) require selecting a specific $k$ for the multi-step horizon, potentially leaving critical control information on the table. Further, it has been shown that multi-step inverse models can be insufficient when the dynamics are periodic (Levine et al., 2024). Action-bisimulation uses a soft invariance metric to extend single-step models, which better preserves long-term controllability.
Bisimulation methods. Bisimulation describes future invariant state representations, originally applied to stationary representations (Larsen & Skou, 1989; Dean et al., 1997; Ferns et al., 2004), before being extended to continuous state MDPs (Ferns et al., 2011). Reward-based bisimulation methods have gained popularity through learned deep representations (Zhang et al., 2020b). This has been extended to non-optimal policies (Castro et al., 2021), with generalized value function bounds (Kemertas & Aumentado-Armstrong, 2021) and augmented with state discretization (Kemertas & Jepson, 2022) and clustering (Liu et al., 2023). Bisimulation-based methods have also been applied in different contexts: expert policy similarity (Agarwal et al., 2021; Bertran et al., 2022; Mazoure et al., 2021), goal-conditioned RL (Hansen-Estruch et al., 2022) and reward-action policy equivalence (Liao et al., 2023; Castro, 2020). While this work draws on reward-bisimulation, action-bisimulation is fundamentally offline and task-agnostic because it takes an expectation over actions, removing its dependence on any policy.
@@ 3 Preliminaries
A Markov decision process is defined by the tuple $\mathcal{M}\coloneqq(\mathcal{S},\mathcal{A},p,R)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space and $s\in\mathcal{S},a\in\mathcal{A}$ are states and actions respectively. $p(s^{\prime}|s,a)$ is the transition function that gives the probability of the next state $s^{\prime}$ given the current state and action $(s,a)$. The reward function $R(s,a)$ maps state and action to a scalar reward. A policy $\pi(a|s)$ is the probability of an action given the current state.
This work utilizes the following two-phase paradigm: in the first phase, the agent first takes actions without access to the reward function $R(s,a)$ to generate a dataset of ordered state action tuples $\mathcal{D}\coloneqq\{(s^{(0)},a^{(0)}),\ldots(s^{(|\mathcal{D}|-1)},a^{(|\mathcal{D}|-1)})\}$. Then, a representation $\phi:\mathcal{S}\rightarrow\mathcal{Z}$ is learned from $\mathcal{S}$. In the second phase, the agent learns from extrinsic reward utilizing the learned representation.
@@
$d_{\text{r-bisim}}(s_{i},s_{j})=\max_{a}\underbrace{(1-c)\cdot|R(s_{i},a)-R(s_{j},a)|}_{\text{base case}}+\underbrace{cW_{1}(d)(p(\cdot|s_{i},a),p(\cdot|s_{j},a))}_{\text{recursive step}},$ (3)
where $W_{1}(d)$ is the 1-Wasserstein distance and $c$ is a scalar hyperparameter that weights the multi-step sensitivity of the distance. The 1-Wasserstein metric measures the distance between next-state distributions in the latent bisimulation space. We propose a novel controllability-based relation, which replaces reward equivalence with single-step control equivalence. By replacing rewards in the equivalence, the relation is task-agnostic.
@@
$d_{\text{a-bisim}}(s_{i},s_{j},\psi)=\underbrace{(1-c)\cdot\|\psi(s_{i})-\psi(s_{j})\|_{1}}_{\text{base case}}+\underbrace{c\cdot\mathbb{E}_{a\sim U(\mathcal{A})}\left[W_{1}(p(\cdot|s_{i},a),p(\cdot|s_{j},a))\right]}_{\text{recursive step}}$ (6)
@@ 4.2 Action-Bisimulation Metric
In the off-policy setting, we propose using one of two expectations for the recursive step: over the uniform distribution of actions $E_{a\sim U(\mathcal{A})}$ or over the behavior distribution: $E_{a\sim\pi_{b}(s_{i})}$. The use of the behavioral distribution applies to settings where random actions might restrict the distribution of observed states. In practice, these are computed using the empirical mean. Then, the action-controllability bisimulation metric using learned models is:
$d_{\text{a-bisim}}(s_{i},s_{j},\psi_{\theta},\phi_{\eta})=(1-c)\cdot\|\psi_{\theta}(s_{i})-\psi_{\theta}(s_{j})\|_{1}+c\cdot\mathbb{E}_{a\sim U(\mathcal{A})}\left[W_{1}(f(\phi_{\eta}(s_{i}),a),f(\phi_{\eta}(s_{j}),a))\right]$ (9)
@@
$L(\mathcal{D})=\frac{1}{N}\sum_{s_{i},s_{j}\sim\mathcal{D}}\left|\|\phi_{\eta}(s_{i})-\phi_{\eta}(s_{j})\|_{1}-d_{\text{a-bisim}}(s_{i},s_{j},\psi,\phi)\right|.$ (10)
@@
  Input: Dataset without reward $(s,a,s^{\prime})\sim\mathcal{D}$, initial encoder $\phi^{\bar{\theta}}(s)$
@@ 5 Experiments
We evaluate experiments in three domains illustrated in Figure LABEL:fig:downstream_rl. Nav2D is a 15x15 grid environment where the agent navigates using cardinal directions to the center of the grid, avoiding randomly generated 2x2 obstacles. Pointmaze (Fu et al., 2020) is a 2D Mujoco control environment where the agent takes actions to reach a goal location while being impeded by obstacles. We also investigate Distractor Pointmaze where the background in Pointmaze has been replaced with photorealistic distractions in the form of video clips. Finally, Habitat (Savva et al., 2019b) is a complex 3D environment where the agent must navigate through scans of human environments to reach a goal location. Additional environment details are in Appendix H (number of obstacles, goal/grid size, randomization, etc.) and all other relevant hyperparameters are in Appendix I.
@@ 5.2 Downstream Learning
As we can see in Figure LABEL:fig:downstream_rl, learning with the action-bisimulation representation outperforms other methods in terms of sample efficiency, even k-step controllability (ACRO), by a substantial margin. This provides evidence for hypotheses 1 and 2, that action-bisimulation learns useful representations which compare well with other methods. That reconstruction and data augmentation-based methods bVAE and CURL perform poorly is not unexpected: in this domain, the agent is often small, so these methods achieve low reconstruction loss even when they omit the most important element: the agent position. On the other hand, SSI captures the agent position but is highly myopic, limiting transfer to downstream tasks. We hypothesize ACRO struggles because it relies on predicting action from two states separated by $k$ timesteps, which is ill-posed, especially when using a dataset of random actions. Additional details on baselines can be found in Appendix G.
@@ 5.3 Background Distractors
In this section we evaluate hypothesis 3: whether the action-bisimulation encoding is robust to distractors. We assess this through a modified Pointmass environment with a photorealistic visual background. The foreground, that is the agent, goal position, and obstacles, remain the same. We visualize the distractor environments in Figure LABEL:paired_images, where the agent has been exaggerated.
@@ 6 Conclusion
Controllability-capturing encodings for reinforcement learning are a promising direction for representation pretraining since they can be learned without reward but are still able to filter out uncontrollable distractors. However, existing methods either only capture short-term controllability or are dependent on demonstration data, which has implicit task bias. We introduce the action-bisimulation encoding, which builds off of myopic representations by enforcing recursive invariance to learn a supervision-free multi-step controllability representation. The empirical results in this work demonstrate how these encodings can be used to improve the sample efficiency, especially in domains with significant background distractors. The primary limitation of this method is the inverse dynamics single-step model, which might not capture all controllable features, just a subset. This can result in the representation being agnostic to important task elements. A more in-depth discussion of limitations is included in Appendix F. Altogether, action-bisimulation is a novel invariance relation for capturing controllability from offline data that removes expert performance requirements and smoothly handles long-horizon controllability.
7 Acknowledgements
This work has taken place in part in the Safe, Correct, and Aligned Learning and Robotics Lab (SCALAR) at The University of Massachusetts Amherst. SCALAR research is supported in part by the NSF (IIS-2323384), AFOSR (FA9550-20-1-0077), and the Center for AI Safety (CAIS). The work was supported by the National Defense Science & Engineering Graduate (NDSEG) Fellowship sponsored by the Air Force Office of Science and Research (AFOSR). Special thanks to collaborators Stephen Guigere, Harshit Sikchi, Alex Levine, Siddhant Agarwal, Rudolf Lioutikov, Yuchen Cui, Akanksha Saran, Wonjoon Goo, Daniel Brown, Prasoon Goyal, Christina Yuan, and Ajinkya Jain for their fruitful conversations and timely help.
@@ C.2 K-step Base Cases for Action-Bisimulation
C.3 Adaptive Regularization for Minimal Representation
To train an encoder with the loss described in Eq. 7, it is necessary to choose a regularizing constant $\beta$ beforehand. We found it was possible (and sometimes easier from a hyper-parameter search perspective) to adapt the $\beta$ parameter to the current performance of encoder $\psi_{\theta}$. We changed $\beta$ throughout training according to the accuracy of the inverse dynamics predictions, lowering the regularization constant when the accruacy was low and raising it when the accuracy was high. The intuition is that if accuracy is low, then the representation needs to be less minimal and so we need to regularize less heavily. We calculated the regularization constant $\beta_{i}$ where $i$ is the training iteration with:
$\beta_{i}=\beta_{\text{max}}(1-\exp(-4\alpha_{i-1}^{2})),$
where $\beta_{\text{max}}$ is the maximum regularization constant and $\alpha_{i-1}$ is the action prediction accuracy during the previous iteration. This trick did not significantly impact our results, but lessened the hyper-parameter search.
@@ Appendix E Additional RL Results
As we can see, Figure 6 provides evidence that action-bisimulation encodings are not limited to discrete actions.
@@
Appendix F Limitations
The three primary limitations we observed in this work for implementing action-bisimulation are as follows, and we go into further detail in the subsequent subsections:
  1. 1.
    The minimum controllable single-step representation, especially when using learned inverse dynamics can omit controllable information if the action is overrepresented in the state
  2. 2.
    Uncontrollable, but reward-relevant elements must be incorporated into the representation after it is learned.
  3. 3.
    Both the forward model for transitions and the action-bisimulation encoding representation are bootstrapped over the expectation over all actions. This can result in unstable training.
  4. 4.
    When a task does not require much lookahead, action-bisimulation will only provide a marginal benefit.
F.1 Limitations of Minimum Controllable Representation
@@
For example, consider the scenario where in the top corner of the screen there is a small display of the last action that the agent takes. In this scenario, the inverse dynamics model is likely to learn to only pay attention to this part of the screen, ignoring other components such as the state of the agent. This is because paying attention to this part is a sufficient representation of actions, even though it does not capture all action-relevant information. Information-based methods such as Equation 16 or generative controllability representations are a possible solution for this, but we have found empirically that they do not seem to learn representations useful for RL. As a result, future work is for investigating action-controllable components that are not necessary for inverse dynamics (action prediction).
F.2 Uncontrollable Reward-relevant components
Another possible limitation is relevant for all controllability pertaining: a representation that captures controllable elements may fail to capture uncontrollable reward-relevant components. For example, consider a goal-based task where the goal is part of the state. The goal itself is not controllable and thus the representation of the goal will not be encoded in an action-bisimulation representation. While the action-bisimulation method is task-agnostic, at least insofar as the initial offline dataset is task-agnostic, it is not a sufficient representation in every task.
This issue can be mitigated by a variety of approaches. The simplest is to simply allow the representation to be modified to be task-specific in the RL training, and we employ this strategy in this work. However, more complex strategies might add an additional channel for task-relevant information, or integrating classic reward-bisimulation to learn the task-relevant components on top of the pre-trained action-bisimulation ones.
F.3 Training Instability of Bootsrapping
One of the challenges when learning an action-bisimulation representation is the inherent bootstrapping where the forward model is trained with $f(\phi(s),a)\rightarrow\phi^{\prime}(s)$, and the encodings themselves are being updated with Equation 10. Since the action distribution in Equation 9 is over the uniform expectation over actions, this can result in instability because of the combinatorial complexity of actions. One way we mitigated this is through the adaptive learning rate of the forward model, but future work should investigate stabilizing the convergence, especially if action-bisimulation is applied to online data.
F.4 Tasks without Lookahead
Finally, while multi-step controllability is a powerful property, not all tasks require this kind of lookahead, and it is not clear that multi-step pretraining would outperform single-step or other baselines in these cases. For example, in the popular Mujoco locomotion domains (Todorov et al., 2012), knowing about future control can often be distracting to the agent—all the relevant information is captured by determining how the current action will affect future actions. Domains where long-term control is useful, such as manipulation, can also be challenging for the current form of action-bisimulation because of the minimal sufficient information property of the single-step losses. Future work is aimed at investigating this in greater detail.
@@ I.3 Habitat
PPO Parameters
Parameter Value clip_param 0.2 ppo_epoch 4 num_mini_batch 2 value_loss_coef 0.5 entropy_coef 0.01 lr 0.00025 eps .00001 max_grad_norm 0.5 num_steps 128 hidden_size 512 gamma 0.99 tau 0.95